A silversmith combined pure silver that cost $16.30 per ounce with 50 oz of a silver alloy that cost $15.00 per ounce. How many ounces of pure silver were used to make an alloy of silver costing $15.80 per ounce?

Answer by richwmiller(17219)

16.30x+50*15=15.80*(50+x)
x is number of ounces of pure silver
x=80
check
16.30*80+50*15=15.80*130
16.30*80+50*15=2054
1304+750=2054
2054=2054
ok